# Fourier Circuits in Neural Networks and Transformers: A Case Study of Modular Arithmetic with Multiple Inputs

### v3

@@ 2 RELATED WORK
Algebraic Tasks Learning Mechanism Interpretability. The study of neural networks trained on algebraic tasks has been pivotal in shedding light on their training dynamics and inductive biases. Notable contributions include the work of [69, 31, 71] on modular addition and subsequent follow-up studies, investigations into learning parities [22, 6, 84, 86, 85, 73, 106, 100], and research into algorithmic reasoning capabilities [74, 33, 47, 57, 21, 14, 87, 64, 105, 91, 34]. The field of mechanistic interpretability, focusing on the analysis of internal representations in neural networks, has also seen significant advancements through the works of [15, 67, 61, 72, 95, 20, 78, 46, 17, 54] and others.
@@ 3.1 Data and Network Setup
where with $u_{i,j}(a_{j})$ being the $a_{j}$-th component of $u_{i,j}$. We consider the parameter set: $\Theta:=\{\|\theta\|_{2,k+1}\leq 1\},\text{~{}~{}where~{}}\|\theta\|_{2,k+1}:=(\sum_{i=1}^{m}\|\theta_{i}\|_{2}^{k+1})^{\frac{1}{k+1}},$ and $\|\theta_{i}\|_{2}:=(\sum_{j=1}^{k}\|u_{i,j}\|^{2}_{2}+\|w_{i}\|_{2}^{2})^{\frac{1}{2}}.$
Here $\|\theta\|_{2,k+1}$ is the $L_{2,k+1}$ matrix norm of $\theta$ (Definition D.2), and $\|\theta_{i}\|_{2}$ is the $L_{2}$ vector norm of the concatenated vector of the parameters in $\theta_{i}$. The training objective over $\Theta$ is then as follows.
@@ 3.2 Margins of the Neural Networks
We denote $g:\mathbb{R}^{U}\times\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ as the margin function, where for given $(x,y)\in D_{p}$, $g(\theta,x,y):=f(\theta,x)[y]-\max_{y^{\prime}\in\mathcal{Y}\backslash\{y\}}f(\theta,x)[y^{\prime}].$
@@
The margin for a given dataset $D_{p}$ is denoted as $h:\mathbb{R}^{U}\rightarrow\mathbb{R}$ where $h(\theta):=\min_{(x,y)\in D_{p}}g(\theta,x,y).$
@@
where the first step is from Definition 3.4, the second step is from Definition 3.3 and the last step is from the linearity of the expectation. Now, we introduce an important concept of a duality stationary pair $(\theta^{*},q^{*})$.
@@
We have $g^{\prime}$ uses a weighted sum rather than max, so $g(\theta,x,y)\leq g^{\prime}(\theta,x,y)$. Following linearity of expectation, we get the expected class-weighted margin as
@@ 3.3 Connection between Training and the Maximum Margin Solutions
Figure 1: Cosine shape of the trained embeddings (hidden layer weights) and corresponding power of Fourier spectrum. The two-layer network with $m=2944$ neurons is trained on $k=4$-sum mod-$p=47$ addition dataset. We even split the whole datasets ($p^{k}=47^{4}$ data points) into the training and test datasets. Every row represents a random neuron from the network. The left figure shows the final trained embeddings, with red dots indicating the true weight values, and the pale blue interpolation is achieved by identifying the function that shares the same Fourier spectrum. The right figure shows their Fourier power spectrum. The results in these figures are consistent with our analysis statements in Lemma 4.2. See Figure 57 in Appendix J.2 for similar results when $k$ is 3 or 5.
@@
Therefore, to comprehend the global minimize, we can explore the maximum-margin solution as a surrogate, enabling us to bypass complex analyses in non-convex optimization. Furthermore, [58] states that under the following condition, the maximum-margin solutions and class-weighted maximum-margin ($g^{\prime}$) solutions are equivalent with each other.
@@ 4 MAIN RESULT
We characterize the Fourier features to perform modular addition with $k$ input in the one-hidden-layer neuron network. We show that every neuron only focuses on a distinct Fourier frequency. Additionally, within the network, there is at least one neuron for each frequency. When we consider the uniform class weighting, where $\mathcal{L}_{\lambda}(\theta)$ is based on $\tau(a_{1},\dots,a_{k})[c^{\prime}]:=1/(p-1)~{}~{}\forall c^{\prime}\neq a_{1}+\dots+a_{k},$ we have the following main result:
Theorem 4.1 (Main result, informal version of Theorem I.2).
@@
See formal proof in Appendix I.2. By Lemma 4.2, we get $\gamma^{*}$ and the single-neuron class-weighted maximum-margin solution set $\Omega_{q}^{{}^{\prime}*}$. By satisfying Condition 3.8, we know it is used in the maximum-margin solution. By Lemma 4.3, we can construct the network $\theta^{*}$ that uses neurons in $\Omega_{q}^{{}^{\prime}*}$. By Lemma E.2, we know that it is the maximum-margin solution. Finally, by Lemma H.2, we know that all frequencies are covered. ∎
@@
Figure 2: All Fourier spectrum frequencies being covered and the maximum normalized power of the embeddings (hidden layer weights). The one-hidden layer network with $m=2944$ neurons is trained on $k=4$-sum mod-$p=47$ addition dataset. We denote $\widehat{u}[i]$ as the Fourier transform of $u[i]$. Let $\max_{i}|\widehat{u}[i]|^{2}/(\sum|\widehat{u}[j]|^{2})$ be the maximum normalized power. Mapping each neuron to its maximum normalized power frequency, (a) shows the final frequency distribution of the embeddings. Similar to our construction analysis in Lemma 4.3, we have an almost uniform distribution over all frequencies. (b) shows the maximum normalized power of the neural network with random initialization. (c) shows, in frequency space, the embeddings of the final trained network are one-sparse, i.e., maximum normalized power being almost 1 for all neurons. This is consistent with our max-margin analysis results in Lemma 4.3. See Figure 6 and  8 in Appendix J.2 for results when $k$ is 3 or 5.
@@ 4.1 TECHNIQUE OVERVIEW
In this section, we propose techniques overview of the proof for our main result. We use $\mathbf{i}$ to denote $\sqrt{-1}$. Let $f:\mathbb{Z}_{p}\rightarrow\mathbb{C}$. Then, for each frequency $j\in\mathbb{Z}_{p}$, we define $f$ discrete Fourier transform (DFT) as $\widehat{f}(j):=\sum_{\zeta\in\mathbb{Z}_{p}}f(\zeta)\exp(-2\pi\mathbf{i}\cdot j\zeta/p).$ Let $\Omega_{q}^{{}^{\prime}*}$ be the single neuron class-weighted maximum-margin solution set (formally defined in Definition F.6). We first show how we get $\Omega^{\prime*}_{q}$.
Lemma 4.2 (Informal version of Lemma F.8).
@@
See formal proof in Appendix F.5. The proof establishes the maximum-margin solution’s sparsity in the Fourier domain through several key steps. Initially, by Lemma F.7, focus is directed to maximizing Eq. (F.4). For odd $p$, Eq. (F.4) can be reformulated with magnitudes and phases of $\widehat{u}_{i}$ and $\widehat{w}$ (discrete Fourier transform of $u_{i}$ and $w$), leading to an equation involving cosine of their phase differences. Plancherel’s theorem is then employed to translate the norm constraint to the Fourier domain. This allows for the optimization of the cosine term in the sum, effectively reducing the problem to maximizing the product of magnitudes of $\widehat{u}_{i}$ and $\widehat{w}$ (Eq. (20)). By applying the inequality of arithmetic and geometric means, we have an upper bound for the optimization problem. To achieve the upper bound, equal magnitudes are required for all $\widehat{u}_{i}$ and $\widehat{w}$ at a single frequency, leading to Eq. (24). The neurons are finally expressed in the time domain, demonstrating that they assume a specific cosine form with phase offsets satisfying certain conditions. ∎
@@
Lemma 4.3 (Informal version of Lemma G.3).
@@
See formal proof in Appendix G.3. Our goal is to show that $2^{2k-1}\cdot{p-1\over 2}$ neurons $\theta_{i}^{*}\in\Omega_{q}^{{}^{\prime}*}$ are able to simulate $\frac{p-1}{2}$ type of $\cos$ computation. We have the following expansion function of $\cos_{\zeta}(x)$, which denotes $\cos(2\pi\zeta x/p)$.
@@
The above equation can decompose a $\cos(\sum)$ to some basic elements. We have $2^{k}$ terms in the above equation. By using the following fact in Lemma G.1,
@@
where each term can be constructed by $2^{k-1}$ neurons. Therefore, we need $2^{k-1}2^{k}$ total neurons. To simulate $\frac{p-1}{2}$ type of simulation, we need $2^{2k-1}\frac{p-1}{2}$ neurons. Then, using the Lemma E.1, we construct the network $\theta^{*}$. By using the Lemma E.2 from [58], we get it is the maximum-margin solution. ∎
@@ 5 EXPERIMENTS
First, we conduct simulation experiments to verify our analysis for $k=3,4,5$. Then, we show that the one-layer transformer learns 2-dimensional cosine functions in their attention weights. Finally, we show the grokking phenomenon under different $k$. Please refer to Appendix J.1 for details about implementation.
One-hidden Layer Neural Network. We conduct simulation experiments to verify our analysis. In Figure 1 and Figure 2, we use SGD to train a two-layer network with $m=2944=2^{2k-2}\cdot(p-1)$ neurons, i.e., Eq. (2), on $k=4$-sum mod-$p=47$ addition dataset, i.e., Eq. (1). Figure 1 shows that the networks trained with SGD have single-frequency hidden neurons, which support our analysis in Lemma 4.2. Furthermore, Figure 2 demonstrates that the network will learn all frequencies in the Fourier spectrum which is consistent with our analysis in Lemma 4.3. Together, they verify our main results in Theorem 4.1 and show that the network trained by SGD prefers to learn Fourier-based circuits. There are more similar results when $k$ is 3 or 5 in Appendix J.2.
Figure 3: 2-dimension cosine shape of the trained $W^{KQ}$ (attention weights) and their Fourier power spectrum. The one-layer transformer with attention heads $m=160$ is trained on $k=4$-sum mod-$p=31$ addition dataset. We even split the whole datasets ($p^{k}=31^{4}$ data points) into training and test datasets. Every row represents a random attention head from the transformer. The left figure shows the final trained attention weights being an apparent 2-dim cosine shape. The right figure shows their 2-dim Fourier power spectrum. The results in the figures are consistent with Figure 1. See Figure 9 and Figure 10 in Appendix J.3 for similar results when $k$ is 3 or 5.
One-layer Transformer. We find similar results in one-layer transformers. Let $E$ be input embedding and $W^{P},W^{V},W^{K},W^{Q}$ be projection, value, key and query matrix. The $m$-heads attention layer can be written as
@@
We denote $W^{K}W^{Q\top}$ as $W^{KQ}$ and call it attention matrix. In Figure 3 , we train a one-layer transformer with $m=160$ heads attention and hidden dimension 128, i.e., above equation, on $k=4$-sum mod-$p=31$ addition dataset, i.e., Eq. (1). Figure 3 shows that the SGD trained one-layer transformer learns 2-dim cosine shape attention matrices, which is similar to the one-hidden layer neural networks in Figure 1. This means that the attention layer has a learning mechanism similar to neural networks in the modular arithmetic task. It prefers to learn (2-dim) Fourier-based circuits when trained by SGD. There are more similar results when $k$ is 3 or 5 in Appendix J.3.
@@
Figure 4: Grokking (models abruptly transition from bad generalization to perfect generalization after a large number of training steps) under learning modular addition involving $k=2,3,4,5$ inputs. We train two-layer transformers with $m=160$ attention heads on $k=2,3,4,5$-sum mod-$p=97,31,11,5$ addition dataset with $50$% of the data in the training set under AdamW [48] optimizer 1e-3 learning rate and 1e-3 weight decay. We use different $p$ to guarantee the dataset sizes are roughly equal to each other. The blue curves show training accuracy and the red ones show validation accuracy. There is a grokking phenomenon in all figures. However, as $k$ increases, the grokking phenomenon becomes weak. See explanation in Section 5.

### v4

@@ 1 INTRODUCTION
Detailed comparison with [65]. Theoretically, we generalize beyond the results of [65] to $k$ inputs. There are unique technique challenges for our setting and not presented in previous settings: (1) While constructing a general $k$ version of the max-margin solution, we need our unique sum-to-product Identities for $k$ inputs, which was proved by our Lemma E.1; (2) To calculate the number of neurons, we need our unique Lemma 4.3 and Eq (29) to handle cosine operation on $k$ inputs; (3) We need our Lemma D.7 and Lemma D.8 to handle multiple variable Fourier transform, where we also introduce multiple inequalities for $k$ inputs version. Empirically, our experiments verified that the theoretical insight obtained can be carried over to practical transformers. Furthermore, the study of grokking is beyond [65], as the study of grokking is only possible when there are $k\geq 2$. Our general $k$ version is necessary for studying some key properties of network learning, like grokking over tasks of increasing complexity or generalization ability over tasks of increasing complexity.
@@ 2 RELATED WORK
Algebraic Tasks Learning Mechanism Interpretability. The study of neural networks trained on algebraic tasks has been pivotal in shedding light on their training dynamics and inductive biases. Notable contributions include the work of [76, 32, 78] on modular addition and subsequent follow-up studies, investigations into learning parities [23, 6, 92, 94, 93, 81, 115, 108], and research into algorithmic reasoning capabilities [82, 34, 49, 64, 22, 14, 95, 71, 113, 99, 35]. The field of mechanistic interpretability, focusing on the analysis of internal representations in neural networks, has also seen significant advancements through the works of [15, 74, 68, 79, 103, 21, 86, 47, 17, 59, 80, 48, 18, 60, 54] and others.
@@
Theoretical Work About Fourier Transform. To calculate Fourier transform there are two main methodologies: one uses carefully chosen samples through hashing functions (referenced in works like [38, 37, 44, 45]) to achieve sublinear sample complexity and running time, while the other uses random samples (as discussed in [10, 36, 72]) with sublinear sample complexity but nearly linear running time. There are many other works studying Fourier transform [87, 40, 33, 63, 19, 89, 16, 90, 20, 96, 56].
@@ 3.1 Data and Network Setup
where with $u_{i,j}(a_{j})$ being the $a_{j}$-th component of $u_{i,j}$. We consider the parameter set:
$\Theta:=$$\{\|\theta\|_{2,k+1}\leq 1\},$
[…]
$\|\theta_{i}\|_{2}:=$ $(\sum_{j=1}^{k}\|u_{i,j}\|^{2}_{2}+\|w_{i}\|_{2}^{2})^{\frac{1}{2}}.$
Here $\|\theta\|_{2,k+1}$ is the $L_{2,k+1}$ matrix norm of $\theta$ (Definition B.2), and $\|\theta_{i}\|_{2}$ is the $L_{2}$ vector norm of the concatenated vector of the parameters in $\theta_{i}$. The training objective over $\Theta$ is then as follows.
@@ 3.2 Margins of the Neural Networks
We denote $g:\mathbb{R}^{U}\times\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ as the margin function, where for given $(x,y)\in D_{p}$,
$g(\theta,x,y):=f(\theta,x)[y]-\max_{y^{\prime}\in\mathcal{Y}\backslash\{y\}}f(\theta,x)[y^{\prime}].$
@@
The margin for a given dataset $D_{p}$ is denoted as $h:\mathbb{R}^{U}\rightarrow\mathbb{R}$ where
$h(\theta):=\min_{(x,y)\in D_{p}}g(\theta,x,y).$
@@
where the first step is from Definition 3.4, the second step is from Definition 3.3, and the last step is from the linearity of the expectation. Now, we introduce an important concept of a duality stationary pair $(\theta^{*},q^{*})$.
@@
We have $g^{\prime}$ uses a weighted sum rather than max, so $g(\theta,x,y)\leq g^{\prime}(\theta,x,y)$. Following the linearity of expectation, we get the expected class-weighted margin as
@@ 3.3 Connection between Training and the Maximum Margin Solutions
Figure 1: Cosine shape of the trained embeddings (hidden layer weights) and corresponding power of Fourier spectrum. The two-layer network with $m=2944$ neurons is trained on $k=4$-sum mod-$p=47$ addition dataset. We even split the whole datasets ($p^{k}=47^{4}$ data points) into the training and test datasets. Every row represents a random neuron from the network. The left figure shows the final trained embeddings, with red dots indicating the true weight values, and the pale blue interpolation is achieved by identifying the function that shares the same Fourier spectrum. The right figure shows their Fourier power spectrum. The results in these figures are consistent with our analysis statements in Lemma 4.2. See Figure 57 in Appendix H.2 for similar results when $k$ is 3 or 5.
@@
Therefore, to comprehend the global minimize, we can explore the maximum-margin solution as a surrogate, enabling us to bypass complex analyses in non-convex optimization. Furthermore, [65] states that under the following condition, the maximum-margin solutions and class-weighted maximum-margin ($g^{\prime}$) solutions are equivalent to each other.
@@
We have $g^{\prime}(\theta^{*},x,y)=g(\theta^{*},x,y)$ for all $(x,y)\in\operatorname{spt}(q^{*})$, where $\operatorname{spt}$ is the support. It means:
$\{y^{\prime}\in\mathcal{Y}\backslash\{y\}:\tau(x,y)[y^{\prime}]>0\}\subseteq\underset{y^{\prime}\in\mathcal{Y}\backslash\{y\}}{\arg\max}f(\theta^{*},x)[y^{\prime}].$
@@ 4 MAIN RESULT
We characterize the Fourier features to perform modular addition with $k$ input in the one-hidden-layer neuron network. We show that every neuron only focuses on a distinct Fourier frequency. Additionally, within the network, there is at least one neuron for each frequency. When we consider the uniform class weighting, where $\mathcal{L}_{\lambda}(\theta)$ is based on
$\tau(a_{1},\dots,a_{k})[c^{\prime}]:=1/(p-1)~{}~{}\forall c^{\prime}\neq a_{1}+\dots+a_{k},$ (5)
we have the following main result:
Theorem 4.1 (Main result, informal version of Theorem G.2).
@@
See formal proof in Appendix G.2. By Lemma 4.2, we get $\gamma^{*}$ and the single-neuron class-weighted maximum-margin solution set $\Omega_{q}^{{}^{\prime}*}$. By satisfying Condition 3.8, we know it is used in the maximum-margin solution. By Lemma 4.3, we can construct the network $\theta^{*}$ that uses neurons in $\Omega_{q}^{{}^{\prime}*}$. By Lemma C.2, we know that it is the maximum-margin solution. Finally, by Lemma F.2, we know that all frequencies are covered. ∎
@@
Our analysis essentially provides hints on how neural networks learn to perform well. Note that humans do modular calculations completely differently with Fourier circuits. Thus, for more general tasks, the model behavior may differ from that of human beings. On the other hand, the Fourier spectrum feature pattern could be useful for out-of-distribution (OOD), robust learning, or designing better learning algorithms, e.g., making the implicit regularization explicit.
Figure 2: All Fourier spectrum frequencies being covered and the maximum normalized power of the embeddings (hidden layer weights). The one-hidden layer network with $m=2944$ neurons is trained on $k=4$-sum mod-$p=47$ addition dataset. We denote $\widehat{u}[i]$ as the Fourier transform of $u[i]$. Let $\max_{i}|\widehat{u}[i]|^{2}/(\sum|\widehat{u}[j]|^{2})$ be the maximum normalized power. Mapping each neuron to its maximum normalized power frequency, (a) shows the final frequency distribution of the embeddings. Similar to our construction analysis in Lemma 4.3, we have an almost uniform distribution over all frequencies. (b) shows the maximum normalized power of the neural network with random initialization. (c) shows, in frequency space, the embeddings of the final trained network are one-sparse, i.e., maximum normalized power being almost 1 for all neurons. This is consistent with our max-margin analysis results in Lemma 4.3. See Figure 6 and  8 in Appendix H.2 for results when $k$ is 3 or 5.
@@ 4.1 Technique Overview
In this section, we propose techniques overview of the proof for our main result. We use $\mathbf{i}$ to denote $\sqrt{-1}$. Let $f:\mathbb{Z}_{p}\rightarrow\mathbb{C}$. Then, for each frequency $j\in\mathbb{Z}_{p}$, we define $f$ discrete Fourier transform (DFT) as $\widehat{f}(j):=\sum_{\zeta\in\mathbb{Z}_{p}}f(\zeta)\exp(-2\pi\mathbf{i}\cdot j\zeta/p).$ Let $\Omega_{q}^{{}^{\prime}*}$ be the single neuron class-weighted maximum-margin solution set (formally defined in Definition D.6).
First, we provide a brief informal high-level intuition of our proof. Based on Condition 3.8 (proved by Lemma 4.3), the maximum-margin solutions and class-weighted maximum-margin are equivalent to each other. Thus, in Lemma 4.2, we can get the class-weighted maximum-margin solution as the problem has been reduced from a network max-margin problem to a single-neuron max-margin problem. Finally, combining all of these, we have our main Theorem 4.1.
Now, we show how to get $\Omega^{\prime*}_{q}$.
Lemma 4.2 (Informal version of Lemma D.8).
@@
See formal proof in Appendix D.5. The proof establishes the maximum-margin solution’s sparsity in the Fourier domain through several key steps. Initially, by Lemma D.7, focus is directed to maximizing Eq. (D.4). For odd $p$, Eq. (D.4) can be reformulated with magnitudes and phases of $\widehat{u}_{i}$ and $\widehat{w}$ (discrete Fourier transform of $u_{i}$ and $w$), leading to an equation involving cosine of their phase differences. Plancherel’s theorem is then employed to translate the norm constraint to the Fourier domain. This allows for the optimization of the cosine term in the sum, effectively reducing the problem to maximizing the product of magnitudes of $\widehat{u}_{i}$ and $\widehat{w}$ (Eq. (21)). By applying the inequality of arithmetic and geometric means, we have an upper bound for the optimization problem. To achieve the upper bound, equal magnitudes are required for all $\widehat{u}_{i}$ and $\widehat{w}$ at a single frequency, leading to Eq. (25). The neurons are finally expressed in the time domain, demonstrating that they assume a specific cosine form with phase offsets satisfying certain conditions. ∎
@@
Lemma 4.3 (Informal version of Lemma E.3).
@@
See formal proof in Appendix E.3. Our goal is to show that $2^{2k-1}\cdot{p-1\over 2}$ neurons $\theta_{i}^{*}\in\Omega_{q}^{{}^{\prime}*}$ are able to simulate $\frac{p-1}{2}$ type of $\cos$ computation. We have the following expansion function of $\cos_{\zeta}(x)$, which denotes $\cos(2\pi\zeta x/p)$.
@@
The above equation can decompose a $\cos(\sum)$ to some basic elements. We have $2^{k}$ terms in the above equation. By using the following fact in Lemma E.1,
@@
where each term can be constructed by $2^{k-1}$ neurons. Therefore, we need $2^{k-1}2^{k}$ total neurons. To simulate $\frac{p-1}{2}$ type of simulation, we need $2^{2k-1}\frac{p-1}{2}$ neurons. Then, using the Lemma C.1, we construct the network $\theta^{*}$. By using the Lemma C.2 from [65], we get it is the maximum-margin solution. ∎
@@ 5 EXPERIMENTS
First, we conduct simulation experiments to verify our analysis for $k=3,4,5$. Then, we show that the one-layer transformer learns 2-dimensional cosine functions in their attention weights. Finally, we show the grokking phenomenon under different $k$. Please refer to Appendix H.1 for details about implementation.
5.1 One-hidden Layer Neural Network
We conduct simulation experiments to verify our analysis. In Figure 1 and Figure 2, we use SGD to train a two-layer network with $m=2944=2^{2k-2}\cdot(p-1)$ neurons, i.e., Eq. (2), on $k=4$-sum mod-$p=47$ addition dataset, i.e., Eq. (1). Figure 1 shows that the networks trained with SGD have single-frequency hidden neurons, which support our analysis in Lemma 4.2. Furthermore, Figure 2 demonstrates that the network will learn all frequencies in the Fourier spectrum, which is consistent with our analysis in Lemma 4.3. Together, they verify our main results in Theorem 4.1 and show that the network trained by SGD prefers to learn Fourier-based circuits. There are more similar results when $k$ is 3 or 5 in Appendix H.2.
Figure 3: 2-dimension cosine shape of the trained $W^{KQ}$ (attention weights) and their Fourier power spectrum. The one-layer transformer with attention heads $m=160$ is trained on $k=4$-sum mod-$p=31$ addition dataset. We even split the whole datasets ($p^{k}=31^{4}$ data points) into training and test datasets. Every row represents a random attention head from the transformer. The left figure shows the final trained attention weights being an apparent 2-dim cosine shape. The right figure shows their 2-dim Fourier power spectrum. The results in the figures are consistent with Figure 1. See Figure 9 and Figure 10 in Appendix H.3 for similar results when $k$ is 3 or 5.
5.2 One-layer Transformer
We find similar results in one-layer transformers. Let $E$ be input embedding and $W^{P},W^{V},W^{K},W^{Q}$ be projection, value, key and query matrix. The $m$-heads attention layer can be written as
@@ 5.2 One-layer Transformer
We denote $W^{K}W^{Q\top}$ as $W^{KQ}$ and call it attention matrix. In Figure 3, we train a one-layer transformer with $m=160$ heads attention and hidden dimension 128, i.e., above equation, on $k=4$-sum mod-$p=31$ addition dataset, i.e., Eq. (1). Figure 3 shows that the SGD-trained one-layer transformer learns 2-dim cosine shape attention matrices, which is similar to the one-hidden layer neural networks in Figure 1. This means that the attention layer has a learning mechanism similar to neural networks in the modular arithmetic task. It prefers to learn (2-dim) Fourier-based circuits when trained by SGD. There are more similar results when $k$ is 3 or 5 in Appendix H.3.
@@
Figure 4: Grokking (models abruptly transition from bad generalization to perfect generalization after a large number of training steps) under learning modular addition involving $k=2,3,4,5$ inputs. We train two-layer transformers with $m=160$ attention heads on $k=2,3,4,5$-sum mod-$p=97,31,11,5$ addition dataset with $50$% of the data in the training set under AdamW [50] optimizer 1e-3 learning rate and 1e-3 weight decay. We use different $p$ to guarantee the dataset sizes are roughly equal to each other. The blue curves show training accuracy, and the red ones show validation accuracy. There is a grokking phenomenon in all figures. However, as $k$ increases, the grokking phenomenon becomes weak. See explanation in Section 5.
5.3 Grokking under Different $k$